### v1

# Assessing differences between local galaxy dust attenuation and point source extinction within the same environments

Abstract
Context. Dust attenuation in galaxies has often been used as a proxy for the extinction of point sources, such as supernovae, even though this approach ignores fundamental differences between the two cases.
Aims. We present an analysis of the impact of geometric effects and scattering within dusty media on recovered galaxy dust properties.
Methods. We employ SKIRT, a radiative transfer code, to simulate observations of point sources embedded in dust clouds, as well as spiral and elliptical galaxies. We examine various galaxy morphologies, inclinations, and instrument apertures.
Results. We find that in galaxies the scattering of light into the line of sight and the presence of sources at different depths within the galaxy make attenuation fundamentally different from extinction. For a medium with intrinsic extinction curve slope $R_{V}=3.068$, we recover effective attenuation curve slopes $R_{V_{eff}}$ ranging from $0.5$ to $7$, showing that the two quantities are not analogous, even for local resolved observations. We find that $R_{V_{eff}}$ greatly depends on dust density, galaxy morphology, and inclination, the latter being the most significant. A single simulated galaxy, viewed from different angles, can qualitatively reproduce the well-known relation between attenuation strength $A_{V_{eff}}$ and $R_{V_{eff}}$ observed for star-forming galaxy samples. An increase in dust density leads to higher $R_{V_{eff}}$ across all inclinations, which, assuming a correlation between stellar mass and dust density, explains the increase in $R_{V_{eff}}$ with mass observed in star-forming galaxies. However, we are unable to explain the differences in $R_{V_{eff}}$ between star-forming (higher $R_{V_{eff}}$) and quiescent (lower $R_{V_{eff}}$) high-mass galaxies.
Conclusions. We conclude that highly attenuated regions of simulated face-on galaxies yield $R_{V_{eff}}$ within a $10\%$ error of the intrinsic extinction curve slope of the medium, allowing for the distinction of different dust types. For edge-on spirals, however, the median $R_{V_{eff}}$ for low $A_{V_{eff}}$ regions appears to better approximate the extinction curve slope.
Key Words.:
dust, extinction - galaxies: ISM - galaxies: structure
1 Introduction
Type Ia supernovae (SNe Ia) have long served as fundamental distance indicators for observational cosmology. They have contributed to the measurement of the accelerated expansion of the Universe (Riess et al., 1998; Perlmutter et al., 1999) and to the constraint of various cosmological parameters, such as the Hubble constant $H_{0}$ (e.g. Dhawan et al., 2018; Riess et al., 2022; Galbany et al., 2023). However, when comparing the values of $H_{0}$ obtained via SNe Ia and other distance indicators of the distance ladder with those obtained from early-time probes, a tension of $\sim 5\sigma$ has been reported (Planck Collaboration et al., 2020; Riess et al., 2020). Although several cosmological model extensions have been proposed as the solution to this tension (e.g. Karwal & Kamionkowski, 2016; Zhao et al., 2017; Raveri, 2020), the possibility that some systematic error is, at least partially, responsible has not yet been fully discarded. In fact, recent results show that by using an inverse distance ladder where SN distances are calibrated via baryon acoustic oscillations (BAO) data, an $H_{0}$ compatible with early-time measurements can be recovered (Camilleri et al., 2025), pointing to the fact that, from the SN Ia point of view, the tension might originate in the standard local distance ladder (Wojtak & Hjorth, 2024).
The process of SN Ia calibration for cosmology relies on a number of empirical corrections based on observed color-luminosity and light-curve shape-luminosity (Phillips, 1993; Tripp, 1998) relations, which have usually been assumed to be universal (Guy et al., 2010). However, SNe Ia originating in high-mass galaxies have been shown to be more luminous than those originating in low-mass galaxies after the shape and color corrections (Sullivan et al., 2010; Kelly et al., 2010; Lampeitl et al., 2010). This effect is known as the “mass step” and it could hint toward the existence of a systematic bias in the calibration, which could also potentially influence SN Ia cosmological results.
One likely source for the mass-step as a systematic in calibrated SN luminosities is the incomplete treatment of the previously mentioned color-luminosity relation. On top of its assumption of universality, it also combines two different physical effects in the same linear relation: i) intrinsic color-luminosity relations, i.e., fainter SNe Ia are redder; and ii) extrinsic dust effects, i.e., more dust extinction leads to fainter magnitudes and redder SNe through selective absorption. González-Gaitán et al. (2021) propose the existence of two separate populations of SNe Ia as an explanation for the mass-step, with each having different intrinsic color-luminosity properties and dust environment characteristics. This is supported by Grayling et al. (2024), who find evidence for both intrinsic and dust environment differences in SN Ia from different host galaxy mass populations. Brout & Scolnic (2021) suggest that host galaxy dust differences are solely responsible for the effect and define two separate $R_{V}$ populations for low- and high-mass galaxies, finding significantly different values for each one and an overall reduction in the mass-step. Similarly, Wiseman et al. (2022) find that using separate $R_{V}$ populations for younger and older galaxies results in a similar improvement, while Johansson et al. (2021) recover no mass-step when using an individual best-fit $R_{V}$ for each SN Ia in the calibration. Duarte et al. (2023) find that SN host galaxy properties can be used to define a “dust-step”, which nevertheless fails to fully account for the mass-step. It has also been shown that apparent galactic dust properties can vary considerably between different galaxies and that they strongly correlate with environmental properties such as the stellar mass and the stellar age of the population in question (e.g., Garn & Best, 2010; Zahid et al., 2013; Salim et al., 2018; Duarte et al., 2023). As such, assuming a standard universal dust effect for SN Ia might prove inadequate in correcting the color-luminosity relation and could ultimately be, at least partially, responsible for the observed mass-step.
In an effort to improve SN Ia calibration, many methodologies have been recently proposed (e.g. Brout & Scolnic, 2021; Wiseman et al., 2022; Duarte et al., 2023), relying on the assumption that galactic dust properties can be used as an adequate proxy for SN Ia extinction. These approaches, while in some cases presenting a degree of success in reducing the Hubble residuals for the calibrated SN, might ultimately prove to be flawed, given that attenuation and extinction are not the same. While they appear similar, the reddening laws describing the effects of dust on SN and other point source observations, also known as extinction laws, are fundamentally different from the reddening laws affecting integrated light observations such as galaxies, also known as attenuation laws (Krügel, 2009; Salim & Narayanan, 2020). This also means that the parameters usually used to describe dust effects, while sharing the same nomenclature in most of the literature, must be interpreted with care and in different ways.
As light travels through the interstellar medium (ISM) and interacts with dust particles, it experiences absorption and scattering effects, which are especially prevalent for blue wavelengths (Calzetti et al., 1994). These effects lead both to a dimming and a reddening of the original spectral energy distribution, which can be parameterized by an empirical reddening law (e.g. Cardelli et al., 1989; Fitzpatrick, 1999; Calzetti, 2001).
For a point source, the total extinction for the $V$ band, $A_{V}$, which serves as a normalization for the extinction curve, describes the dust column density in the optical path of the observed light. Likewise, the color excess for the $B-V$ color index, $E(B-V)$, refers only to the reddening experienced along that single line of sight, which means that the total-to-selective extinction $R_{V}=A_{V}/E(B-V)$ is only a function of the optical properties of the dust grains themselves (Draine, 2011).
For an extended stellar population or a galaxy, this picture is complicated as a result of two primary factors. On the one hand, the necessity of using a larger aperture for galaxy observations makes the light scattered back into the line of sight a relevant component of the overall observed flux, with the strength of this scattered component being dependent on the morphology and inclination of the galaxy, as demonstrated by Chevallard et al. (2013). Those authors show that galaxy inclination can drastically impact attenuation properties, meaning that identical galaxies viewed from different angles can produce wildly different attenuation curve strengths and slopes. On the other hand, the interspersion of stars and dust lanes leads to varying column densities inside the galaxy, along with different stellar light contributions. Viaene et al. (2017) found that the thicker the dust lane, the more the emission is dominated by stars located in the outermost regions of the galaxy, which are not meaningfully attenuated.
When taken together, the above effects mean that, in the case of a galaxy, the values measured for $A_{V_{eff}}$ and $R_{V_{eff}}$ stop being representative of the true optical properties of the medium along the line of sight and should instead be understood as effective values. As such, they cannot be easily related to the extinction experienced by SN light, even if the local dust environment is properly mapped (Duarte et al., 2023).
Although this has been long known, it has only recently permeated the SN field. Chevallard et al. (2013) and Viaene et al. (2017) show that a deeper knowledge of the underlying galaxy geometry is needed to properly describe dust attenuation. Duarte et al. (2023) show that the SN Ia calibration returns much higher Hubble residuals when galactic dust properties are used to directly constrain the color luminosity correction. Popovic et al. (2024) conclude that dust in a SN line of sight cannot be accurately described using its host galaxy dust properties.
In this work we present a comprehensive analysis of how point source extinction relates to galactic attenuation. We rely on radiative transfer simulations to systematically study how factors such as instrument aperture and galaxy inclination, morphology and dust composition influence the recovered values of $A_{V_{eff}}$ and $R_{V_{eff}}$ and how these relate to the real optical properties of the medium. More specifically, we seek to understand under which conditions, if any, a host galaxy dust measurement can be said to truly reflect the extinction experienced by a local point source, such as a SN, and therefore be used as its proxy. We use the radiative transfer code SKIRT (Camps & Baes, 2015, 2020) to simulate observations for both point sources and a variety of galaxies (spiral and elliptical), observed from multiple angles and with differing apertures, the details of which are presented in Section 2. In Section 3, we present and discuss our simulation results and compare them with observational data, focusing on how well purely geometrical effects can explain observed $A_{V_{eff}}-R_{V_{eff}}$ relations. In Section 4, we analyze the link between dust properties and stellar mass, an environmental property related to SN Ia calibration, and to what extent it can be explained by our simulations. We also discuss how dust attenuation data could be used to quantitatively recover true extinction $R_{V}$ values, as well as the impacts of the source spectrum on the recovered dust properties. Finally, in Section 5, we summarize our main conclusions.
2 Methods
In this work, we make use of SKIRT, a powerful radiative transfer code that uses the Monte Carlo method (Camps & Baes, 2015, 2020) [^1] https://skirt.ugent.be/root/_home.html. This code relies on the simulation of the path of a number of photon packets emitted from a light source through a gridded astrophysical dusty medium. By taking into account both absorption and scattering effects, the code is able to stochastically simulate a view of the model from any orientation, as observed by a particular instrument. While it is also possible to simulate the contributions of dust emission, we do not do so, as it is not particularly relevant for the wavelengths in question in this work.
We are interested in simulating the basic geometries of both galaxies and point sources. With this goal, we focus on three highly abstracted models: 1) a point source immersed within a galactic dust cloud; 2) a spiral galaxy-like model; 3) an elliptical galaxy-like model. The point source system is mainly used as an ideal benchmark for comparison with the galaxy models.
The simulated astrophysical structure in a SKIRT model is made up of two systems, a light source and a dust medium system. The spectral energy distribution (SED) for the source system is defined based on the simple stellar populations (SSPs) described by Bruzual & Charlot (2003), assuming a Chabrier (2003) initial mass function (IMF), with a metallicity $Z=0.02$ and an age of $t_{age}=5$ Gyr. [^2] Given that SKIRT simulations output both the simulated observed flux and the total emitted flux, we can directly compute $A_{V_{eff}}$ and $R_{V_{eff}}$, meaning that the recovered dust properties are free from any degeneracies and that they are not impacted by the SED chosen for the source system.. To simulate aperture photometry observations, $10^{8}$ photon packets are used, while to simulate full-galaxy field observations (with a pixel size corresponding to a $25$ pc resolution), $10^{11}$ photon packets are used instead [^3] The value of $10^{11}$ photon packets, despite being quite high, was chosen because, for any lower number, severe stochastic fluctuations in $R_{V_{eff}}$ were present for the majority of the pixels.. We make use of SKIRT’s “ExtinctionOnly” mode given that, as stated above, we are not interested in dust emission.
By default, a Zubko et al. (2004) dust mix model is used for the dust medium, composed of bare silicate, graphite and polycyclic aromatic hydrocarbon (PAH) dust grains, with half of the PAH grains being neutral and half being ionized. A Mathis-Rumpl-Nordsiek (MRN) dust mix, composed of bare silicate and graphite (Mathis et al., 1977), is also used for a specific example, discussed in Section 3.2.2. The simulation outputs both the emitted and the observed flux, meaning that we can obtain accurate values for both $A_{V}$ and $R_{V}$ ($A_{V_{eff}}$ and $R_{V_{eff}}$, if we are in an attenuation scenario). While these two parameters are not ideal for a complete analysis of extinction and attenuation curves, given that they are restricted to the behavior of only two wavelengths, they nevertheless remain one of the most common parametrizations for dust effects, both in SN and galaxies.
We simulate observations from a variety of angles and apertures. We simulate global aperture photometry for an array of aperture SED detectors located at a distance of $10$ Mpc from the center of the model. The detectors are positioned at different angular positions, ranging from a face-on view of the galaxy at $0^{\circ}$ (along the z-axis) to an edge-on view at $90^{\circ}$ (along the xy-plane). The detector aperture radius is varied to examine different observing conditions. We also simulate full-galaxy field data for a detector that records the SED for each pixel in the field of view, also located at a distance of $10$ Mpc from the center of the model. A physical pixel size corresponding to a $25$ pc resolution is used.
2.1 Point Source/Dust Cloud Model
[…]
Based on Baes et al. (2000), the dust cloud system density profile is described by a flattened spheroidal geometry given by Eq. 1:
$\rho(R,z)=\frac{1}{q}\ \rho\left(\sqrt{R^{2}+\frac{z^{2}}{q^{2}}}\right),$ (1)
where $q=0.5$ is the flattening parameter and $\rho$ is the Plummer (1911) density profile, given by Eq. 2:
$\rho(r)=\rho_{0}\left(1+\frac{r^{2}}{c^{2}}\right)^{-5/2},$ (2)
where $c=4000$ pc is the scale length. The normalization $\rho_{0}$ is varied to produce different optical depths in the dust system.
2.2 Spiral Galaxy Model
The light source system for the spiral galaxy is divided into two components: a disk and a bulge. The disk is described by a disk geometry characterized by a double-exponential profile given by Eq. 3:
[…]
The bulge component is described by a flattened spheroidal geometry, defined by Eq. 1 with a flattening parameter of $q=0.7$ and a Sérsic (1963) $\rho(r)$ density profile given by Eq. 4:
$\rho(r)=\rho_{0}\ S_{n}\left(\frac{r}{r_{eff}}\right),$ (4)
[…]
The dust medium geometry is taken to be the same as that of the spiral source disk, described by a disk characterized by a double-exponential profile as defined by Eq. 3, with the normalization $\rho_{0}$ being varied to produce simulated galaxies with different optical depths. We use a scale length parameter $h_{R}=4000$ pc and a scale height parameter $h_{z}=350$ pc. Even though the dust disk is typically found to be thinner than the stellar disk (e.g., De Geyter et al., 2014), we opt for using disks with the same thickness. We do so because we are primarily interested in the interaction between light and dust and thus want to limit the number of stars effectively outside the dust disk. However, we find that using a thinner dust disk does not significantly alter the results discussed below.
2.3 Elliptical Galaxy Model
[…]
Once more following Baes et al. (2000), the dust system for the elliptical galaxy is described by a flattened spheroidal geometry, defined by Eq. 1 with a flattening parameter of $q=0.5$ and a Plummer (1911) $\rho(r)$ density profile, as described by Eq. 2 with $c=4000$ pc. The normalization $\rho_{0}$ is varied to produce different optical depths in the dust system.
3 Results
3.1 Point Source Model
In this section we present an analysis of simulated photometric observations for the point source model. Particularly, we analyze the impact of different apertures on the dust properties recovered for the model and the transition between extinction and attenuation regimes.
The observed flux can be divided into two components: 1) the direct flux, comprising photon packets that reach the detector without interacting with dust; 2) the scattered flux, comprising packets that have gone through at least one scattering event, ending up in the line of sight. Fig. 1 shows the ratio between these two fluxes for face-on observations of a point source inside a galactic dust cloud as seen by simulated instruments with different aperture radii. The dust density profile was normalized so that the total optical depth along the x-axis is $\tau_{V_{X}}=5$.
Figure 1: Ratio between the scattered (into the line of sight) and direct fluxes as a function of the wavelength $\lambda$ for face-on observations of a simulated point source, immersed in a galactic dust cloud. Simulated observations for a variety of aperture radii are shown. The cloud dust density has been normalized so that the total optical depth along the x-axis is $\tau_{V_{X}}=5$.
As expected, as the size of the aperture used for the observations increases, so does the relative contribution of the scattered flux to the total flux. If the aperture is small enough, as in the case with a radius of $1$ pc, the scattered component is virtually non-existent and we are left with only the direct flux component. Clearly, the larger apertures place us in the attenuation regime, meaning that the recovered dust properties should be understood as effective, as they are no longer representative of the actual optical properties of the medium along the line of sight.
This is further illustrated in Fig. 2, which shows, as an example, the values of $A_{V_{eff}}$ and $R_{V_{eff}}$ for a variety of observation angles when using two different simulated instruments, with aperture radii of $1$ AU (left) and $20$ kpc (right), respectively. As hinted before, the observed behavior greatly depends on the aperture used. For the smaller aperture, we recover a constant value of $R_{V}=3.068$ that does not depend on the observation angle and matches exactly the value of the extinction curve slope for the Zubko dust mix $R_{V_{Zubko}}$, as directly computed from the material extinction cross-sections outputted by SKIRT. Likewise, the values of $A_{V}$ can be taken as true indicators of dust extinction, as they correctly reflect the amount of light that has been absorbed or scattered away from the line of sight.
For the larger aperture, we note that the recovered $A_{V_{eff}}$ values are $57\%$ to $43\%$ lower than the respective true extinction $A_{V}$. This is a result of the scattered light that is now contributing to the observations, as seen if Fig. 1. Moreover, we note that the relation between $A_{V}$ and $A_{V_{eff}}$ is close to linear, indicating that within the same dust cloud the scattered and absorbed fluxes scale consistently at this wavelength. Additionally, we find that $R_{V_{eff}}$ is highly dependent on the galaxy inclination, ranging from $R_{V_{eff}}=2.212$, for face-on observations, to $R_{V_{eff}}=3.120$, for edge-on observations. We conclude that, for this particular example, an increase in the observation aperture can result in a $28\%$ deviation from the intrinsic optical extinction $R_{V}$.
Figure 2: Values of effective $R_{V_{eff}}$ as a function of $A_{V_{eff}}$ for a simulated point source immersed in a galactic dust cloud. Different colors denote different observation angles, from $0^{\circ}$ (face-on) to $90^{\circ}$ (edge-on). The results from two different instruments are presented: One with an aperture radius of $1$ AU (left) and another with an aperture radius of $20$ kpc (right). The cloud dust density has been normalized so that the total optical depth along the x-axis is $\tau_{V_{X}}=5$.
Figure 3: Values of effective $R_{V_{eff}}$ as a function of aperture radius for a simulated point source immersed in a galactic dust cloud. Results are shown in blue for $\tau_{V_{X}}=5$, orange for $\tau_{V_{X}}=15$ and green for $\tau_{V_{X}}=25$. Face-on results are represented by dots, while edge-on results are represented by squares. The dashed black line reflects the true $R_{V}$ of the medium.
We can refine our analysis by looking at the $R_{V_{eff}}$ values recovered for each of the apertures discussed in Fig. 1, which are plotted in Fig. 3. In addition to the $\tau_{V_{X}}=5$ normalization discussed above, results for two other dust densities, $\tau_{V_{X}}=15$ and $\tau_{V_{X}}=25$, are presented, for both face-on and edge-on observations. We find that, for the current setup, any aperture with a radius larger than $25-50$ pc results in a $R_{V_{eff}}$ that deviates from $R_{V_{Zubko}}$. In particular, as stated above, we confirm that galaxy inclination has a great impact on both $A_{V_{eff}}$ and $R_{V_{eff}}$, as face-on and edge-on simulated observations differ greatly from each other.
The differences in the simulated $R_{V_{eff}}$ arise from the interplay between the direct and scattered fluxes. In general, the scattered flux tends to dominate the bluer wavelengths, with the direct flux dominating the redder wavelengths, as shown in Fig. 4. As we move from a face-on view to an edge-on view, the scattered flux becomes even bluer when compared to the direct flux. Overall, this results in a flatter edge-on attenuation curve, boasting higher values of $R_{V_{eff}}$ than those observed for the face-on view, as seen in Fig. 3. This effect is primarily driven by the thickening of the optical light path for the edge-on observations, which increases the interaction probability between light and the medium, contributing to the absorption and, most relevant, the scattering of more blue light away from the direct flux. Although this increase in optical depth is also expected to make the scattered flux overall redder, this component still becomes bluer when compared to the direct flux. Likewise, if the dust density is increased, we expect an increase in $R_{V_{eff}}$ across all observation angles, which is precisely what is observed from the simulations. As such, a higher dust column density actually brings the face-on $R_{V_{eff}}$ values closer to $R_{V_{Zubko}}$, while it increases the deviation for the edge-on values.
Figure 4: Flux components for face-on (top) and edge-on (bottom) simulated observations of a point source immersed in a galactic dust cloud with $\tau_{V_{X}}=5$. The observations were simulated for an aperture radius of $r=20$ kpc. The emitted flux, total observed scattered and direct fluxes are shown in purple, blue, orange and green, respectively.
We highlight that the results discussed in this section are independent of the SED chosen for the light-source system and thus reflect the composition and geometry of the dust medium. In fact, repeating the simulations for a light source with $t_{age}=10$ Myr, we recover the same exact values of $A_{V_{eff}}$ and $R_{V_{eff}}$, apart from expected stochastic variations inherent to the method.
We also note that our point source model is not representative of SN observations. While the angular resolution of most ground based SN Ia surveys would place them above our threshold for the extinction regime, the fact is that, due to their short lifetimes, light can only effectively travel and scatter within a small vicinity around the transient (roughly $0.1$ pc in the first four months after explosion). This means that, unless surrounded by a very dense and exotic dust cloud that could give rise to light echoes, as described in Wang et al. (2008) and Bulla et al. (2018), the SN-dust interaction should always be properly described by extinction. These SN light-echoes are usually only visible at very late stages in the SN evolution. However, the scattering effect described above might be important for other long-lived point sources like stars (e.g. Cepheids) or even compact stellar clusters, taking into account that their source luminosity is fainter than a SN.
3.2 Galaxy Models
Having established a baseline with point source observations, in this section we present an analysis of simulated photometric observations for galaxy models. This differs from the previous example in that an extended light-source distribution is used instead of a single point source. It is particularly useful to look at two different observation scenarios: global observations, resulting from aperture photometry on the entire galaxy, and local observations, resulting from a full-galaxy pixel-to-pixel 2D map.
3.2.1 Global Observations
Fig. 5 shows the simulated values of $R_{V_{eff}}$ as a function of $A_{V_{eff}}$ for both a spiral and an elliptical-like galaxy model, observed from a variety of angles with an aperture radius of $20$ kpc. In both models, the dust density profile was normalized so that the total optical depth along the x-axis is $\tau_{V_{X}}=5$.
Figure 5: Values of $R_{V_{eff}}$ as a function of $A_{V_{eff}}$ (top) and the corresponding values $A_{V_{dir}}$ and $R_{V_{dir}}$, calculated from only the direct flux component, ignoring the scattered flux component (bottom). Results for a simulated spiral (left) and elliptical galaxy (right) are shown. Different colors denote different observation angles, from $0^{\circ}$ (face-on) to $90^{\circ}$ (edge-on). An aperture radius of $20$ kpc was used. The galaxies dust density has been normalized so that the total optical depth along the x-axis is $\tau_{V_{X}}=5$. The dashed black line reflects the true extinction $R_{V_{Zubko}}$.
Generally, for both galaxy types, $R_{V_{eff}}$ follows a trend similar to that observed for the largest aperture in Fig. 2. However, the two morphologies of the galaxies result in noticeable differences, both in terms of the shape of the $A_{V_{eff}}$-$R_{V_{eff}}$ trend and the $R_{V_{eff}}$ value ranges. In particular, for the spiral galaxy, we find that values can range from $R_{V_{eff}}=0.41$ for face-on observations [^4] The points with $R_{V_{eff}}<1$ are most likely unphysical simulation artifacts, stemming from extremely low values of $A_{V_{eff}}$., to $R_{V_{eff}}=4.86$ for edge-on observations. Likewise, we find that for the elliptical galaxy values range from $R_{V_{eff}}=2.71$, for face-on observations, to $R_{V_{eff}}=5.01$, for edge-on observations. In addition, the observation angle for which $R_{V_{Zubko}}$ is recovered varies greatly between the two, being $\sim 70^{\circ}$ for the spiral and $\sim 30^{\circ}$ for the elliptical.
As mentioned in Section 3.1, the trend between $A_{V_{eff}}$ and $R_{V_{eff}}$ is primarily driven by the scattering of light into the line of sight. This scattered component, when compared to the direct flux, is redder for the face-on observations, lowering the observed $R_{V_{eff}}$, and bluer for the edge-on view, resulting in higher $R_{V_{eff}}$. In addition to this, the fact that the sources now extend to different depths and contribute to the same optical path means that $R_{V_{eff}}$ deviates even more from $R_{V_{Zubko}}$.
We can better understand this light source distribution effect by looking at the attenuation properties that we would recover if only contributions from the direct flux were taken into account. As such, for the same two galaxies analyzed in Fig. 5, we plot the values of $A_{V_{eff}}$ and $R_{V_{eff}}$ that one obtains if the scattered flux is ignored and only the direct flux is taken into account, which we denote as $A_{V_{dir}}$ and $R_{V_{dir}}$, respectively. These results are shown in Fig. 5, from which it becomes clear that, without the reddening effect of the scattered component to counterbalance its effects, the direct flux actually results in higher values of $R_{V_{dir}}$.
From the above, we conclude that the reddening caused by changes in the direct optical paths behaves similarly to the reddening caused by excess scattered light, with $R_{V_{eff}}$ increasing with the observation angle. In this case, this increase is due to the fact that, as we move from a face-on to an edge-on orientation, there is an increase in the number of different optical paths in our observation, as we are probing much deeper into the dust cloud. This also explains why the spiral galaxy face-on direct flux observations in Fig. 5 produce the closest results to $R_{V_{Zubko}}$, as this is the scenario in which the sources are most concentrated, given the narrower depth of the galaxy. We also find that, as observed for the point source, the $A_{V_{eff}}$ values obtained for only the direct flux are systematically higher than those obtained from the full observed flux, as expected.
The effects of considering different dust densities on the effective dust properties are presented in Fig. 6. Similarly to the results shown in Fig. 3, we see that the amount of dust in a galaxy has a great impact on the attenuation properties, namely by increasing the values of $R_{V_{eff}}$ across all observation angles, similarly to what was discussed in Section 3.1. For the normalizations considered, we find that, for spiral galaxies, the values of $R_{V_{eff}}$ range from $0.41$ to $1.54$, at their lowest, and from $4.86$ to $7.38$ at their highest. Likewise, for elliptical galaxies, the values of $R_{V_{eff}}$ range from $2.26$ to $3.31$, at their lowest, and from $3.24$ to $6.16$ at their highest.
This dust density effect is, however, still secondary to the variation of the observation angle, which produces a much larger change in $R_{V_{eff}}$. As such, the effective dust properties of a galaxy appear to be more a function of geometry and morphology than of the dust abundance itself. In practice, this means that, when using global galaxy aperture photometry, the true optical properties of the dust medium, namely $R_{V_{Zubko}}$, become obfuscated and cannot be easily recovered.
Figure 6: Values of $R_{V_{eff}}$ as a function of the total optical depth normalization along the x-axis $\tau_{V_{X}}$, for a spiral (left) and elliptical galaxy (right). Different colors denote different observation angles, from $0^{\circ}$ (face-on) to $90^{\circ}$ (edge-on). An aperture radius of $20$ kpc was used. The dashed black line reflects the true extinction $R_{V_{Zubko}}$.
In Fig. 7, we present a comparison between the $A_{V_{eff}}$-$R_{V_{eff}}$ trend obtained from the various simulated observations detailed in Fig. 6 and the trend obtained for Dark Energy Survey (DES; Abbott et al., 2018) global galaxy fits, as discussed in Duarte et al. (2023). The plot covers different galaxy types, dust densities, and orientations. We find that, with the exception of the absolute $A_{V_{eff}}$ values, our simulations are able to fully reproduce the observed behavior. The mismatch in attenuation values might actually stem from a bias in the DES fits of Duarte et al. (2023), possibly stemming from the Prospector Python package (Johnson et al., 2021). In fact, Zahid et al. (2013) and Salim et al. (2018) recover $A_{V_{eff}}$ value ranges for observed galaxies that are similar to the one obtained from our simulations. The increased level of scatter exhibited by the real data is most likely due to the fact that a real sample is expected to reflect a wide mixture of different morphologies and dust types.
Our results are also consistent with simulated observations of galaxies with different inclinations performed by Chevallard et al. (2013), as well as with other global galaxy observations (e.g Salmon et al., 2016; Leja et al., 2017; Narayanan et al., 2018). Overall, we feel confident in asserting that the observed $A_{V_{eff}}$-$R_{V_{eff}}$ relation is indeed driven by the observation angle, with factors such as dust density and galaxy type contributing mainly to an increase in the trend dispersion.
Figure 7: Values of $R_{V_{eff}}$ as a function of $A_{V_{eff}}$ for the simulated galaxies presented in Fig. 6 (blue circles, lower axis) and DES global galaxy fits from Duarte et al. (2023) (orange squares, upper axis). The plot illustrates the same general trend, even though the $A_{V_{eff}}$ values are not compatible.
3.2.2 Local Observations
[…]
Contrary to what we found for a point source, observing with a $25$ pc resolution appears to be insufficient to meaningfully mitigate the scattered flux component across the entire galaxy, which now comes from integrated light. Even so, it appears that, for regions such as the centers of the galaxies, where the dust density is high, the scattered light is indeed negligible. In fact, looking at Fig. 9, where maps of $A_{V_{eff}}$ and $R_{V_{eff}}$ for these two galaxies are shown, we see that, for these central regions, we recover a value of $R_{V_{eff}}$ close to $R_{V_{Zubko}}$. As we move away from the center of the galaxy, we find a decrease in the values of $R_{V_{eff}}$. This is consistent with the gradient recovered by Hutton et al. (2015), who find that the observed $R_{V_{eff}}$ values decrease with the galactocentric distance, when using both ultraviolet and optical photometry. While the authors interpret this as evidence for a radial variation of intrinsic dust properties in galaxies, Fig. 9 shows that the same behavior can be recovered with a uniform dust type as a consequence of different levels of scattering and dust-to-stars distributions. There are also some outlier pixels with uncharacteristically high $R_{V_{eff}}$ values, but these arise mainly due to stochastic variations in the simulation and are particularly relevant for thin dust layers, such as the one present in the spiral edge-on view.
Figure 9: Spatial maps of the values of $A_{V_{eff}}$ and $R_{V_{eff}}$, for a simulated spiral (left) and elliptical face-on galaxy (right). A pixel size corresponding to a $25$ pc resolution was used. The galaxies dust density has been normalized so that the total optical depth along the x-axis is $\tau_{V_{X}}=5$. Only pixels with $A_{V_{eff}}>0$ and $R_{V_{eff}}>0$ are shown.
Figure 10: 2D maps of the values for $A_{V_{eff}}$ and $R_{V_{eff}}$ for a face-on simulated spiral (left) and elliptical galaxy (right). The galaxies dust density has been normalized so that the total optical depth along the x-axis is $\tau_{V_{X}}=5$. Each point corresponds to a single pixel with a resolution of $25$ pc, color-coded according to their galactocentric distance. Binned medians are shown in red, with error-bars defined by the bin standard deviation. The dashed black line reflects the true $R_{V}$ of the medium, with the dotted black lines defining a 10% deviation from this value. Only pixels with $A_{V_{eff}}>0$ and $R_{V_{eff}}>0$ are shown.
The relation between the two dust parameters becomes even more apparent if we look at the $A_{V_{eff}}-R_{V_{eff}}$ trends for both face-on galaxy models, plotted in Fig. 10. As stated above, in addition to the light scattered into the line of sight, which tends to be reddened and contributes to drive $R_{V_{eff}}$ down, one must also consider the existence of multiple sources at different depths in the same optical path, which, when taken together, contribute to an increase in $R_{V_{eff}}$. For higher values of $A_{V_{eff}}$, and thus higher dust densities, the amount of light that can escape from the innermost parts of the galaxy, be it scattered or otherwise, drastically goes down. As such, the observed flux becomes largely dominated by the direct flux of stars located in the outermost layers of the galaxy, which mitigates both of the above-mentioned effects.
For this reason, we can recover a value of $R_{V_{eff}}$ within $10\%$ of $R_{V_{Zubko}}$ when observing areas with large optical depth near the center of the galaxies. This recovery is also aided by the high central source density, which also contributes to make the scattered component subdominant in this particular region. We end up slightly overestimating the true extinction value mainly because it is impossible to completely eliminate the contributions from sources deeper in the galaxy, which slightly increased the values of $R_{V_{eff}}$, as seen for the stellar population modeled here.
Looking at the edge-on case for the same two models, plotted in Fig. 11, we see a very similar picture, especially for the elliptical galaxy. The $R_{V_{eff}}$ values are overall higher, as expected from an increase in dust cloud depth and attested by Fig. 6, but converge toward $R_{V_{Zubko}}$ for the highest $A_{V_{eff}}$ regions. The spiral galaxy case presents a few more difficulties, as the $R_{V_{eff}}$ trend never converges to within $10\%$ of $R_{V_{Zubko}}$. The most likely explanation for this is that light from the deepest layers of the galaxy is not being sufficiently attenuated for the outermost stars to dominate the emission, leading to $R_{V_{eff}}$ values much higher than $R_{V_{Zubko}}$, even for regions with higher $A_{V_{eff}}$. In fact, it appears that, despite the high level of dispersion, the low $A_{V_{eff}}$ regions are much better at tracing $R_{V_{Zubko}}$, possibly because these are also the regions with the fewest stars. Overall, it appears that, to ensure that the $R_{V_{eff}}$ values recovered for a galaxy are indeed reflective of $R_{V_{Zubko}}$, we must be sure that both the scattered flux component and the direct component coming from deeper regions of the galaxy are negligible.
Figure 11: Similar to Fig. 10 for simulated edge-on spiral (left) and elliptical galaxy (right).
From the individual examples discussed above, however, it is not completely clear whether this apparent recovery of $R_{V_{Zubko}}$ for high $A_{V_{eff}}$ regions is to be generally expected or whether it is an artifact of the interplay between these specific dust and source configurations. Thus, we further investigate this effect by examining two slightly modified examples of the elliptical face-on simulation: i) an elliptical galaxy with a lower central source density, which can be achieved using a Sérsic index $n=1$; ii) an elliptical galaxy with a MRN dust mix, instead of a Zubko dust mix, which has an intrinsic extinction $R_{V_{MRN}}=3.65$. The results for the $A_{V_{eff}}-R_{V_{eff}}$ trends, are displayed in Figs. 12 and 13, respectively.
Figure 12: Similar to Fig. 10 for simulated face-on elliptical galaxy with a source distribution given by a Sérsic profile with $n=1$.
Figure 13: Similar to Fig. 10 for a simulated face-on elliptical galaxy with a MRN dust medium.
Despite a much higher level of scatter, Fig. 12 shows that, for higher values of $A_{V_{eff}}$, $R_{V_{eff}}$ continues to approach $R_{V_{Zubko}}$ for the lower central density source distribution. Likewise, using a MRN dust mix substantially changes the value toward which the $R_{V_{eff}}$ trend converges, which is once again very similar to the extinction curve slope value for this particular mix, $R_{V_{MRN}}$.
From these examples we conclude that, to recover a median $R_{V_{eff}}$ within a $10\%$ deviation of the extinction $R_{V}$ in a face-on galaxy, one must focus on the more attenuated areas of a galaxy. The $A_{V_{eff}}$ value for which this threshold is met can range from $40\%$ of the maximum galaxy $A_{V_{eff}}$, in the case of the elliptical galaxy with MRN dust, to $65\%$, in the case of the spiral galaxy. We highlight that these values were derived from a small number of example galaxies and should be taken as merely indicative and that they can vary with observation angle.
One other thing that should be considered is that a single galaxy might host different dust types. In this case, and assuming that the different types are confined to different spatial regions, it is not correct to take the globally most attenuated region as a proxy for the extinction behavior of the whole galaxy. Rather, the most attenuated portions in each region must be considered, with the caveat that it is possible that these will not be dense enough to get within a $10\%$ deviation of the extinction $R_{V}$. Additionally, we must be certain that the interplay between sources and dust in the regions in question is such that the scattered flux is negligible.
We can examine how well these results line up with observations by examining some preliminary dust attenuation values obtained for the All-weather MUse Supernova Integral Field Nearby Galaxies (AMUSING; Galbany et al., 2016) data for the NGC 1448 galaxy. This integral field spectroscopy (IFS) data was matched to GALEX (Martin et al., 2005), HST, 2MASS (Skrutskie et al., 2006) and WISE (Wright et al., 2010) photometry using the piXedfit Python package (Abdurro’uf et al., 2021). The Prospector Python package (Johnson et al., 2021) was used to fit each pixel in a subsection of NGC 1448, with the attenuation properties being expressed in terms of the effective optical depth $\tau_{V_{eff}}$ and the dust index $n$, which controls the slope of the Kriek & Conroy (2013) attenuation curve.
The results for the fitted $\tau_{V_{eff}}$-$n$ trends are displayed in Fig. 14. While these fits are still preliminary, it appears that the same trend observed in Fig. 10, 12 and 13 is present in the AMUSING data, even if expressed through different parameters. It is not totally clear whether the $n$ values are converging toward a value representative of the extinction properties of the medium, or if they are simply clustering near the edge of the prior distribution. Furthermore, our simulations and fits match the results obtained by Decleir et al. (2019), who find a similar increase of the attenuation curve slope with $A_{V_{eff}}$ for resolved observations of NGC 628. These results give some confidence to the applicability of the inferred simulated trends to real galaxies.
Figure 14: Dust index $n$ as a function of the optical depth $\tau_{V_{eff}}$ for fits of a subsection of NGC 1448, using AMUSING IFS data matched with GALEX, HST, 2MASS and WISE photometry. Values for individual pixels are shown in blue, while binned median values are shown in orange.
4 Discussion
4.1 Dust Attenuation and Stellar Mass
Given the nature of the input of the SKIRT simulations, where both the source and dust medium systems require their normalizations to be specified as inputs, a direct probing of the relation between galaxy mass and dust cannot be achieved, as the relation between the two is always determined by the user. However, some important inferences can be drawn from the comparison of our simulations with literature results.
Among others, Garn & Best (2010), Zahid et al. (2013), Salim et al. (2018), Galbany et al. (2022) and Duarte et al. (2023) show that, for star-forming galaxies, there is an increase of $A_{V_{eff}}$ with the stellar mass. It is fair to assume that this is related to an increase in the dust density in high-mass galaxies. Salim et al. (2018) and Duarte et al. (2023) also show that high-mass star-forming galaxies exhibit larger values of $R_{V_{eff}}$. From Fig. 6 it is clear that an increase in the dust density in a galaxy leads directly to an increase in the median values of both $A_{V_{eff}}$ and $R_{V_{eff}}$. Assuming that there is indeed a link between the mass of a galaxy and its dust density, our simulations are able to fully explain the observed increase of $R_{V_{eff}}$ with stellar mass. Crucially, they do so without the need to consider the differing dust types or source SEDs.
Another important result discussed in Zahid et al. (2013), Salim et al. (2018) and Duarte et al. (2023) is that elliptical galaxies are, in general, expected to have much lower values of $A_{V_{eff}}$ and $R_{V_{eff}}$ than high-mass star-forming galaxies. It is this fact that makes it possible for the entire high-mass population to actually have a lower median $R_{V_{eff}}$ than the one observed for the low-mass galaxies. However, this is something our simulations cannot reproduce. If one assumes that the median $A_{V_{eff}}$ observed for a population of elliptical galaxies should be roughly the same as the one observed for a population of low-mass star-forming galaxies, which is corroborated by Zahid et al. (2013), Salim et al. (2018) and Duarte et al. (2023), we can use this assumption to properly normalize the dust contents of both galaxy types. However, under these conditions, we invariably find that the ellipticals boast median $R_{V_{eff}}$ values higher than even those of the high-mass star-forming galaxies. There are two likely explanations. The first is that the simplistic geometries assumed in this work are not able to properly model these two types of galaxies in relation to one another. The second is that there is some fundamental difference between the two galaxy types that goes beyond geometry, such as a different dust composition with different optical properties.
As suggested in Figs. 10, 12 and 13 and tentatively confirmed by Fig. 14, the analysis of full-galaxy field data for the regions of a face-on galaxy with higher $A_{V_{eff}}$ might offer an observational hint toward the true extinction curve slope $R_{V}$ of a given dust sample. While the robustness of this methodology in determining accurate $R_{V}$ values for real galactic dust mediums cannot be ascertained from our simple simulations, observational results are promising, as they show a similar increase in $R_{V_{eff}}$ with $A_{V_{eff}}$ as the one recovered for the simulations. Thus, this might be a very apt method to distinguish between different dust types, perhaps even within the same galaxy. As such, it might be able to provide a reliable way to relate galaxy attenuation data with the extinction of their hosted point sources, such as SNe. This would be invaluable in an effort to better constrain SN Ia calibration and the respective color-luminosity correction.
4.2 Source SED Impact
As mentioned above, when it comes to dust, the results presented in this paper are largely independent of the SED of the simulated sources, given that our simulations output precise fluxes for the emitted and observed fluxes. Because of this, we can avoid many of the uncertainties and degeneracies that are inherent in observations and SED fitting. However, there are a few points to be highlighted.
One of the reasons why our results can be said to be SED independent is that, for any given model simulation, a single SED is used. This means that any effect of the source is factored out when the effective dust properties are computed. We do not, however, expect real galaxies to exhibit a SED that is uniform across all sources, i.e., we do not expect stellar populations of the same age everywhere. In this scenario, both $A_{V_{eff}}$ and $R_{V_{eff}}$ are also expected to be affected by the interplay of the various SEDs, particularly if sources of very different ages are not uniformly distributed within the galaxy and are subject to different levels of attenuation. This will contribute to further deviations from the medium extinction values. Likewise, while the overall luminosity of a galaxy is not a relevant component in our results, given that only the source distribution affects the simulated dust properties, it might play a role on the observed dust properties. In particular, if any of the flux components become dimmer than the limiting magnitude of the observational setup in question, we might end up recovering dust properties that are biased toward the most luminous regions.
5 Conclusions
In this work we have used radiative transfer simulations to explore how SN Ia dust extinction and host galaxy dust attenuation relate to one another. Our main findings are the following:
  1. (i)
    SN Ia extinction (or “point source extinction” as used throughout this paper) and host galaxy attenuation are fundamentally different phenomena. When observing a galaxy, the light scattered into the line of sight becomes a very relevant component of the total observed flux, lowering the value of $R_{V_{eff}}$. In addition, the fact that sources at different depths are contained within the same line of sight contributes to an increase in $R_{V_{eff}}$. When taken in conjunction, these two effects mean that the $R_{V_{eff}}$ measured for any given galaxy is fundamentally different from the optical extinction $R_{V}$ of its particular dust medium, even if small resolutions of $25$ pc are considered.
  2. (ii)
    Global attenuation properties are primarily driven by observation angle. The global $A_{V_{eff}}-R_{V_{eff}}$ trends for observed galaxies can be qualitatively reproduced by a single simulated galaxy observed from different angles. Dust density and galaxy type serve simply to increase the scatter of this trend. We conclude that geometry, rather than dust density or composition, is the primary factor driving global attenuation.
  3. (iii)
    An increase in galaxy dust density results in an increase of $R_{V_{eff}}$ across all observation angles. As such, assuming a correlation between stellar mass and dust density, the geometry of the observations is also able to explain the increase of $R_{V_{eff}}$ values with stellar mass for observed star-forming spiral galaxies.
  4. (iv)
    When spiral and elliptical galaxies are analyzed together, our simulations fail to explain some of the observed behaviors. The literature suggests that ellipticals have $A_{V_{eff}}$ values similar to low-mass spirals. However, if this is assumed in our model, the recovered $R_{V_{eff}}$ values, which are also supposed to be in line with those of low-mass spirals, are much higher than expected. In particular, when taking all galaxies into account, we find that the high-mass population boasts a higher median $R_{V_{eff}}$ than its low-mass counterpart, which is the opposite of the observed trend. This could indicate either a failure in the modeling of one or both galaxy types or that there is something intrinsically different between the two galaxy types beyond geometry, namely their dust properties. In this regard, part of the origin of the mass-step correction in SN Ia distances could indeed be related to differing dust properties among galaxies.
  5. (v)
    The use of the full-galaxy field data opens up the possibility to quantifiably link host galaxy attenuation with point source (such as SNe Ia) extinction. While it is difficult to extract accurate information from individual locations, an analysis of the $A_{V_{eff}}-R_{V_{eff}}$ trend for a galaxy could allow one to infer an approximate value for the extinction $R_{V}$ of the medium and, most importantly, to distinguish between different dust types/compositions. This is tentatively confirmed by preliminary observed galaxy fits. Our examples suggest that, to recover $R_{V_{eff}}$ values within a $10\%$ error of the intrinsic extinction $R_{V}$, areas with high $A_{V_{eff}}$ should be considered, in the case of face-on galaxies. For edge-on spirals, however, the median $R_{V_{eff}}$ for the low $A_{V_{eff}}$ regions – although with a lot of scatter – appears to be better at tracing the intrinsic extinction $R_{V}$. In any case, it is crucial to ensure that these regions have properties that allow for the elimination of both the scattered flux component and the direct component coming from deeper regions of the galaxy.
Acknowledgements.
J. D., S. G. and A. M. acknowledge support by FCT for CENTRA through the Project No. UIDB/00099/2020. J. D. also acknowledges support by FCT under the PhD grant 2023.01333.BD. This work was funded by ANID, Millennium Science Initiative, ICN12_009. Based in part on observations collected at the European Southern Observatory under ESO programme(s) 097.D-0408(A). Based in part on observations made with the NASA/ESA Hubble Space Telescope, obtained at the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS5-26555, associated with programs #17070 and #17191. This publication makes use of data products from the Two Micron All Sky Survey, which is a joint project of the University of Massachusetts and the Infrared Processing and Analysis Center/California Institute of Technology, funded by the National Aeronautics and Space Administration and the National Science Foundation. This work used public data from the GALEX Survey. GALEX is operated for NASA by the California Institute of Technology under NASA contract NAS5-98034. This publication makes use of data products from the Wide-field Infrared Survey Explorer, which is a joint project of the University of California, Los Angeles, and the Jet Propulsion Laboratory/California Institute of Technology, funded by the National Aeronautics and Space Administration.